# Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate

### v1

@@ Abstract
Though free-floating planets (FFPs) that have been ejected from their natal star systems may outpopulate their bound counterparts in the terrestrial-mass range, they remain one of the least explored exoplanet demographics. Due to their negligible electromagnetic emission at all wavelengths, the only observational technique able to detect these worlds is gravitational microlensing. Microlensing by terrestrial-mass FFPs induces rare, short-duration magnifications of background stars, requiring high-cadence, wide-field surveys to detect these events. The Transiting Exoplanet Survey Satellite (TESS), though designed to detect close-bound exoplanets via the transit technique, boasts a cadence as short as 200 seconds and has monitored hundreds of millions of stars, making it well-suited to search for short-duration microlensing events as well. We have used existing data products from the TESS Quick-Look Pipeline (QLP) to perform a preliminary search for FFP microlensing candidates in 1.3 million light curves from TESS Sector 61. We find one compelling candidate associated with TIC-107150013, a source star at $d_{s}=3.194$ kpc. The event has a duration $t_{E}=0.074^{+0.002}_{-0.002}$ days and shows prominent finite-source features ($\rho=4.55^{+0.08}_{-0.07}$), making it consistent with an FFP in the terrestrial-mass range. This exciting result indicates that our ongoing search through all TESS sectors has the opportunity to shed new light on this enigmatic population of worlds.
keywords:
planets and satellites: detection – gravitational lensing: micro – planets and satellites: terrestrial planets
pubyear: 2024pagerange: Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass CandidateSearching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate
@@ 1 Introduction
Free-floating planets (FFPs), planets not bound to any star, constitute an enigmatic planetary demographic of which little is known. Theory and simulation suggest that such planets should be ubiquitous in the Galaxy (Strigari et al., 2012; Hong et al., 2018), with different formation mechanisms dominating at different FFP masses. At high masses ($\gtrsim M_{\text{Jup}}$), they may form in isolation from the collapse of gas, constituting the extreme low-mass end of the stellar mass function (Padoan & Nordlund, 2002; Bonnell et al., 2008; Zwart & Hochart, 2024). FFPs formed through these processes may retain sufficient heat to be detected in the infrared, and recent observations have discovered an unexpectedly large abundance of such worlds (McCaughrean & Pearson, 2023; Pearson & McCaughrean, 2023). At terrestrial masses, however, the dominant formation mechanism is thought to be gravitational ejection from a parent system during the chaotic early phases of system formation, e.g. through planet-planet scattering (Rasio & Ford, 1996; Weidenschilling & Marzari, 1996; Veras & Raymond, 2012), ejection by an inner binary (Nelson, 2003; Sutherland & Fabrycky, 2016; Smullen et al., 2016; Coleman et al., 2023; Standing et al., 2023; Chen et al., 2023; Coleman, 2024), or stellar fly-bys (Wang et al., 2024; Yu & Lai, 2024).
@@
Such terrestrial-mass FFPs likely outpopulate their bound counterparts (Sumi et al., 2023a; Mroz & Poleski, 2023), comprising the majority of Galactic exoplanets in this mass range. However, their negligible emission across the electromagnetic spectrum poses a significant observational challenge. The only technique sensitive to detecting these worlds is gravitational microlensing, in which the gravitational field of the FFP (“lens”) focuses light from a distant background star (“source”), resulting in the temporary apparent magnification of the star as the lens traverses the source along the line of sight. Ground-based microlensing searches have recently yielded the first three observations of terrestrial-mass FFPs (Mróz et al., 2019; Koshimoto et al., 2023; Mróz et al., 2020). Though these observations have confirmed the existence of this population and suggested a high Galactic abundance, ground-based observations are limited by their photometric sensitivity and cadence, with atmospheric interference and technical requirements requiring magnifications of $\gtrsim 10\%$ and a cadence of $>15$ min (e.g., Mróz et al., 2019). Typical microlensing events for terrestrial-mass FFPs have timescales on the order of an hour, and often result in magnifications well below the threshold required by ground-based efforts, making their detection extremely challenging for existing ground-based observatories.
@@
The Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) has provided a relatively unexplored dataset from which new FFPs can be discovered. While TESS was optimized for detecting exoplanets via the transit method, its wide-field imaging capabilities and high-cadence, high-precision photometric observations make it well-suited to also detecting short-duration microlensing events. TESS observes large $24^{\circ}\times 96^{\circ}$ sectors of the sky for 27.4 days at a time, allowing it to monitor millions of stars in its Full-Frame Images (FFIs). TESS’s high photometric precision allows it to detect transit depths (and lensing peaks) as small as a fraction of a percent. Furthermore, its current 200-second FFI cadence enables sensitivity to events lasting significantly less than an hour.
In this work, we leveraged TESS’s unique sensitivity to terrestrial-mass microlensing events to perform a preliminary search for FFPs within TESS Sector 61. We used a dataset consisting of 1.3 million stars with magnitudes as dim as $T=15$ mag observed at a 200-second cadence. This initial search resulted in one potential short-duration microlensing event consistent with a terrestrial-mass FFP.
@@ 2.1 Microlensing Fundamentals
When a foreground FFP crosses near the line of sight to a background source star, its gravitational field perturbs the emitted light rays and creates multiple images of the source. When these multiple images can not be individually resolved, the net effect is an apparent time-varying magnification of the source known as gravitational microlensing. The characteristic gravitational extent of the lens, the Einstein radius, is given by
$R_{E}=\sqrt{\frac{4GMd_{l}(1-d_{l}/d_{s})}{c^{2}}},$ (1)
@@
$t_{E}=\frac{R_{E}}{v_{\text{rel}}},$ (2)
where $v_{\text{rel}}$ is the transverse velocity of the source-lens system relative to the observer. The net magnification of a point-source during a microlensing event can be described analytically by an overall multiplicative factor relative to its baseline flux via (Paczynski, 1986)
@@
When the angular extent of the source $\theta_{s}$ is comparable to or greater than the Einstein angle $\theta_{E}\equiv R_{E}/d_{l}$, the additional parameter $\rho=\theta_{s}/\theta_{E}$ describes the extent of finite-source effects in the light-curve. Finite-source effects generically reduce the peak magnification and elongate the duration of the event (see, e.g., Lee et al. (2009)), resulting in a box-like that makes the light-curve amenable to initial flagging through methods such as the Box Least Squares (BLS) search algorithm (see §3.2 and §4.2).
@@ 2.2 Expected Event Yield
Microlensing is an intrinsically rare phenomenon. Microlensing surveys rely upon the observation of a large number of source stars at a rapid cadence over a long observational baseline. Though TESS’s transit survey was not designed with microlensing in mind, over the course of its Primary and Extended Missions, it has observed hundreds of millions of stars with continuous baselines ranging from 27.4 days to 1 year at cadences as rapid as 200 seconds. As a result, its existing observations are well-suited to a microlensing search for low-mass free-floating planets.
One can get a sense for the expected yield by performing a simple estimate of the event rate for a single source and monochromatic mass distribution of lenses. The rate can be estimated as $\Gamma\approx nv_{\text{rel}}\times(u_{T}R_{E}d_{s})$ (Gaudi et al., 2008), where $n$ is the number density of lenses and $u_{T}R_{E}d_{s}$ is the cross-sectional area spanned by the “microlensing tube” (Griest, 1991), with $u_{T}$ defined as the threshold impact parameter required to produce a detectable magnification, and is $\approx 5$ for TESS’s photometric sensitivity  [^1] One can roughly estimate $u_{T}$ by determining the maximum impact parameter that results in an event satisfying our initial detection threshold (see §3.2), namely a magnification of $A/\sigma_{\text{phot}}>10$ for at least 5 consecutive observations averaged over TESS source stars. $\sigma_{\text{phot}}$ is the median photometric noise from Kunimoto et al. (2022a) scaled to an event duration of $5\times 30\,\text{min}$, where 30 min is the cadence of the TESS Primary Mission. See DeRocco et al. (2024) for further details on estimating $u_{T}$.. Using this expression, a rough estimate of the expected number of detectable low-mass FFP events in a search of existing TESS data is
$N\approx 1~{}\text{event}~{}\Big{(}\frac{u_{T}}{5}\Big{)}\Big{(}\frac{n}{20~{}\rm{pc}^{-3}}\Big{)}\Big{(}\frac{M}{0.1M_{\oplus}}\Big{)}^{1/2}\\ \Big{(}\frac{v_{\text{rel}}}{55~{}\rm{km/s}}\Big{)}\Big{(}\frac{d_{s}}{3~{}\rm{kpc}}\Big{)}^{3/2}\Big{(}\frac{N_{\star}}{6.67\times 10^{7}}\Big{)}\Big{(}\frac{t_{\rm{obs}}}{27.4~{}\rm{days}}\Big{)},$ (5)
where the number of individual single-sector light curves $N_{\star}$ has been chosen to correspond to the TESS dataset of stars down to $T=13.5$ mag for Sectors 1 - 70. We assume a typical source distance $d_{s}=3$ kpc and a per-sector observational baseline $t_{\text{obs}}=27.4$ days.
The value of $n$ has been taken to coincide with the best fit of the FFP mass function suggested by the MOA collaboration (Sumi et al., 2023b) (see also Gould et al. (2022); Mroz & Poleski (2023)) and corresponds to a normalization of $\approx 22$ free-floating planets per star with masses within a half-dex of $M_{\oplus}$. The true abundance of FFPs in this mass range is unknown, however, and may very well be significantly higher. TESS, with its rapid cadence and high photometric sensitivity, is uniquely suited to probe this low-mass region of the mass function. The conservative estimate of Eq. (5) therefore provides considerable motivation to conduct a search for overlooked low-mass FFP microlensing events in TESS observations.
@@ 3.1 Light Curves
We used FFI light curves extracted by the Quick-Look Pipeline (QLP; Huang et al., 2020), which covers all stars in the TESS Input Catalog (currently TICv8.2; Paegert et al., 2021) brighter than $T=13.5$ mag. Since Sector 41, QLP also produces light curves for M dwarfs as faint as $T=15$ mag (Kunimoto et al., 2022b). For TESS Sector 61 (2023 January 18 – February 12), we obtained light curves with 200-second cadence for a total of 1,286,238 targets.
From each light curve, we removed all data points flagged as poor quality as indicated by a quality flag of 1, as well as all data within 0.2 days of the start and end of each TESS orbit, resulting in an average observational baseline per target of 23.6 days. These times correspond to high magnification systematics caused by scattered light from the Earth or Moon. To remove low-frequency trends caused by stellar variability, we applied the biweight detrending algorithm implemented in wotan (Hippke et al., 2019) using a detrending window of 2 days. Given that we expect most terrestrial FFP events to have typical durations of less than one day, the 2-day window was chosen to minimize the impact of event distortion caused by detrending while removing long-term astrophysical trends.
@@ 3.2 Microlensing Event Detection
Transiting exoplanets are commonly found using the BLS period search algorithm (Kovács et al., 2002), which searches for periodic dimming of a target star. The shape of this signal can be approximated as a box, i.e. the dimming initiates and terminates instantaneously. BLS can also be adapted for microlensing studies by searching for non-repeating, brightening events. Though the magnification curve associated with microlensing of point-like sources is not well-approximated by a box, finite-source effects broaden this peak, resulting in light-curves that more closely resemble a box-like plateau (see §4.2 for an example displaying these effects, or Fig. 4 of Johnson et al. (2022) for further discussion). Given that the majority of our TESS targets lie within $\sim 3$ kpc of Earth, terrestrial-mass FFPs are likely to produce events with significant finite-source effects, making the use of the BLS algorithm a reasonable initial means to flag high-SNR events of interest. BLS is also computational inexpensive, which is necessary for searching the large TESS datasets.
@@
$\sigma_{r}^{2}=\sigma_{\mathrm{bin}}^{2}-\sigma_{\mathrm{bin,thy}}^{2},$ (8)
where $\sigma_{\mathrm{bin}}$ is the weighted standard deviation of the residual light curve after binning in time with a bin-size equal to the duration of the transit, and $\sigma_{\mathrm{bin,thy}}$ is the expected standard deviation of the binned light curve if the noise were uncorrelated in time. If $\sigma_{r}$ is estimated to be less than zero, we set $\sigma_{r}=0$.
In order for a signal to be considered an initial candidate to be vetted further, we required $\mathrm{SNR}>10$ and $n\geq 5$.

### v2

# Searching for Free-Floating Planets with TESS: Results from Sectors 61 – 65

@@ Abstract
Though free-floating planets (FFPs) may outpopulate their bound counterparts in the terrestrial-mass range, they remain one of the least explored exoplanet demographics. Due to their negligible electromagnetic emission at all wavelengths, the only observational technique able to detect these worlds is gravitational microlensing. Microlensing by terrestrial-mass FFPs induces rare, short-duration magnifications of background stars, requiring high-cadence, wide-field surveys to detect these events. The Transiting Exoplanet Survey Satellite (TESS), though designed to detect close-bound exoplanets via transits, boasts a Full-Frame Image cadence as short as 200 seconds and has monitored hundreds of millions of stars, providing a unique dataset in which to search for rare short-duration transients. We have performed a preliminary search for FFP microlensing in 7.5 million light curves from TESS Sectors 61–65. We find one short-duration event with a light curve morphology consistent with expectations for a low-mass FFP, but in tension with the expected FFP abundance in this mass range. We consider possible false positive interpretations of this event such as stellar flares, hearbeat binaries, and centrifugal breakout. We find that all interpretations pose some challenges, and discuss the possibility that the event may constitute a first example of a new class of pernicious false positives that future space-based microlensing efforts will encounter. Our ongoing search through the TESS dataset will significantly support the upcoming hunt for rogue worlds with dedicated space-based microlensing surveys, and our results may be used alongside these surveys to place interesting constraints on the spatial distribution of FFPs in the Galaxy.
facilities: Gaia, TESSsoftware: lmfit (Newville et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (Team, 2020; Wes McKinney, 2010), pyLIMA (Bachelet et al., 2017), scipy (Virtanen et al., 2020), tess-point (Burke et al., 2020)
@@ 1 Introduction
Free-floating planets (FFPs), planets not bound to any star, constitute an enigmatic planetary demographic of which little is known. Theory and simulation suggest that such planets should be ubiquitous in the Galaxy (Strigari et al., 2012; Hong et al., 2018), with different formation mechanisms dominating at different FFP masses. At high masses ($\gtrsim M_{\text{Jup}}$), they may form in isolation from the collapse of gas, constituting the extreme low-mass end of the stellar mass function (Padoan & Nordlund, 2002; Bonnell et al., 2008; Portegies Zwart & Hochart, 2024). FFPs formed through these processes may retain sufficient heat to be detected in the infrared, and recent observations have discovered an unexpectedly large abundance of such worlds (McCaughrean & Pearson, 2023; Pearson & McCaughrean, 2023). At terrestrial masses, however, the dominant formation mechanism is thought to be gravitational ejection from a parent system during the chaotic early phases of system formation, e.g. through planet-planet scattering (Rasio & Ford, 1996; Weidenschilling & Marzari, 1996; Veras & Raymond, 2012), ejection by an inner binary (Nelson, 2003; Sutherland & Fabrycky, 2016; Smullen et al., 2016; Coleman et al., 2024; Standing et al., 2023; Chen et al., 2024; Coleman, 2024), or stellar fly-bys (Wang et al., 2024; Yu & Lai, 2024).
@@
Such terrestrial-mass FFPs likely outpopulate their bound counterparts (Sumi et al., 2023; Mroz & Poleski, 2023), comprising the majority of Galactic exoplanets in this mass range. However, their negligible emission across the electromagnetic spectrum poses a significant observational challenge. The only technique sensitive to detecting these worlds is gravitational microlensing, in which the gravitational field of the FFP (“lens”) focuses light from a distant background star (“source”), resulting in the temporary apparent magnification of the star as the lens traverses the source along the line of sight. Ground-based microlensing searches have recently yielded the first three observations of terrestrial-mass FFPs (Mróz et al., 2019; Koshimoto et al., 2023; Mróz et al., 2020). Though these observations have confirmed the existence of this population and suggested a high Galactic abundance, ground-based observations are limited by their photometric sensitivity and cadence, with atmospheric interference and technical requirements requiring magnifications of $\gtrsim 10\%$ and a cadence of >15>15 min (e.g., Mróz et al., 2019). Typical microlensing events for terrestrial-mass FFPs have timescales on the order of an hour, and can have relatively low peak magnifications, making their detection challenging for existing ground-based observatories.
@@
Multiple dedicated space-based microlensing surveys are set to launch in the coming years, opening up a new era in the search for free-floating planets. However, with the increase in sensitivity that these missions will achieve brings new challenges. Chief among them is mitigating potential new classes of false positive that have not been well-characterized by ground-based efforts. The Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) has provided a relatively unexplored dataset from which new short duration signals can be discovered. While TESS was optimized for detecting exoplanets via the transit method, its wide-field imaging capabilities and high-cadence, high-precision photometric observations make it sensitive to detecting brightening features similar to those expected for short-duration microlensing events as well. TESS observes large $24^{\circ}\times 96^{\circ}$ sectors of the sky for 27.4 days at a time, allowing it to monitor millions of stars in its Full-Frame Images (FFIs). TESS’s high photometric precision allows it to detect flux changes as small as a fraction of a percent. Furthermore, its current 200-second FFI cadence enables sensitivity to events lasting significantly less than an hour. While the proximity of most TESS target stars to the Sun results in a low a priori probability of any individual star being lensed by a foreground object over the duration of TESS observations, the large number of stars monitored by TESS over many years, combined with the potentially large population of terrestrial-mass FFP planet in our galaxy inferred from ground-based microlensing surveys, implies that the probability of detecting one or more FFP microlensing event in the extant TESS dataset is non-negligible, as we show Sec. 2.2 and was also shown by Yang et al. (2024).
In this work, we leveraged a subset of the large catalog of existing TESS observations to perform a preliminary search for FFPs within TESS Sectors 61 – 65. We used a dataset consisting of 7.5 million stars with magnitudes as dim as $T=15$ mag observed at a 200-second cadence. This initial search resulted in one potential short-duration microlensing event; however, though the light curve of this event is consistent with that of a terrestrial-mass FFP, we show that the event rate implied by this detection is in tension with expectations for this mass range, making it an unlikely scenario. We therefore examine alternative interpretations of the event, such as a flare, and find that these non-FFP interpretations are also unlikely; however, the limited observational data for such flares leaves us unable to make a quantitative statement about the nature of the event. This underscores the need for potential false positives to be better characterized prior to future space-based microlensing surveys such as the Nancy Grace Roman Space Telescope (Roman; Akeson et al., 2019; Penny et al., 2019) and Earth 2.0 (Ge et al., 2022). This is a task that existing TESS observations are well-suited for.
@@ 2.1 Microlensing Fundamentals
When a foreground FFP crosses near the line of sight to a background source star, its gravitational field perturbs the emitted light rays and creates multiple images of the source. When these multiple images can not be individually resolved, the net effect is an apparent time-varying magnification of the source known as gravitational microlensing. The characteristic angular size of the lens, the Einstein angle, is given by
$\theta_{E}=\sqrt{\frac{4GM(1-d_{l}/d_{s})}{d_{l}c^{2}}},$ (1)
@@
$t_{E}=\frac{\theta_{E}}{\mu_{\text{rel}}},$ (2)
where $\mu_{\text{rel}}$ is the proper motion of the source-lens system relative to the observer. The net magnification of a point-source during a microlensing event can be described analytically by an overall multiplicative factor relative to its baseline flux via (Paczynski, 1986)
@@
When the angular extent of the source $\theta_{s}$ is comparable to or greater than $\theta_{E}$, the additional parameter $\rho=\theta_{s}/\theta_{E}$ describes the extent of finite-source effects in the light-curve. Finite-source effects generically reduce the peak magnification and elongate the duration of the event (see, e.g., Lee et al., 2009), resulting in a box-like morphology that makes the light-curve amenable to initial flagging through methods such as the Box Least Squares (BLS) search algorithm (see §3.2 and §4.2).
@@ 2.2 Expected Event Yield
Microlensing is an intrinsically rare phenomenon. Microlensing surveys rely upon the observation of a large number of source stars at a rapid cadence over a long observational baseline. TESS’s transit survey was not designed with microlensing in mind; its ability to detect such events is limited with respect to dedicated microlensing campaigns due to its short per-sector observational baselines, relatively nearby stars, and choice of target fields. However, over the course of its Primary and Extended Missions, it has observed hundreds of millions of stars with continuous baselines ranging from 27.4 days to 1 year at cadences as rapid as 200 seconds. Though its targets are generally much closer and do not lie in the crowded stellar fields favored by dedicated microlensing searches, TESS’s observations still provide the opportunity to search for short-duration events that may point to a large abundance of Galactic FFPs. Due to TESS’s short observational baseline, any putative detection would need to be cross-referenced with long-baseline ground-based observations in order to better characterize such a signal. Beyond characterizing short-duration false positives for upcoming missions such as Roman and Earth 2.0, there also remains a small possibility of identifying an event that, coupled with other observations, could be identified as an FFP. As such, we performed estimates of the detection rate for such signals. As stated previously, TESS is not designed as a microlensing survey, and therefore the potential detection rate does not compete with those of upcoming missions such as Roman and Earth 2.0; however, as we will show below, even in the case of a null observation in the full TESS dataset, TESS’s existing observations have the potential to help place interesting constraints on the largely uncertain abundance of FFPs in the sub-terrestrial range. Such a constraint would allow more refined predictions of the expected number of detections that Roman and Earth 2.0 may yield, providing further motivation for our ongoing search. Furthermore, by virtue of the proximity of TESS’s targets, TESS data could provide a measurement of or constraint on local FFP abundance, which would prove valuable in characterizing the Galactic distribution of FFPs observed by upcoming microlensing surveys.
We estimate the yield of FFPs in TESS data following the methodology used by Yang et al. (2024). Our expected yields are shown in Figure 1, where each bin displays the expected number of detectable FFPs per dex in mass after searching 67 million TESS QLP light curves ($T<13.5~\mathrm{mag}$) across Sectors 1 – 71. The yields were computed based on assuming a signal-to-noise ratio (SNR) threshold of SNR >10>10 and requiring at least five in-event datapoints. Note that the definition of SNR we employ (which we motivate in §3.2) differs from that of Yang et al. (2024), and results in marginally higher yield estimates than theirs. We adopt the FFP mass function derived by Sumi et al. (2023) using data from existing ground-based observations, and display both the yield assuming the best-fit values for the model as well as the yields for the one standard deviation uncertainties on the fit parameters.
At the best-fit values, we predict $\approx 1$ FFP within the TESS dataset, consistent with the prediction of Yang et al. (2024), while at the $1\sigma$ level we predict as many as $\approx 20$ FFPs. These predictions highlight the present uncertainty on the overall abundance of FFPs in the sub-terrestrial mass range and motivates our ongoing work with TESS, which, even in the case of a null detection, would place improved constraints on the largely-uncertain low-mass FFP abundance.
Figure 1: Expected FFP detection rate per dex in mass, across all currently available TESS sectors from QLP (Sectors 1 – 71). The range of values correspond to different underlying mass functions, based on the central best-fit values describing the Sumi et al. power law and the $\pm 1\sigma$ uncertainties (Sumi et al., 2023). TESS is most sensitive to FFP events between $0.01<M/M_{\oplus}<0.1$.
@@ 3.1 Light Curves
We used FFI light curves extracted by the Quick-Look Pipeline (QLP; Huang et al., 2020), which covers all stars in the TESS Input Catalog (currently TICv8.2; Paegert et al., 2021) brighter than $T=13.5$ mag. Since Sector 41, QLP also produces light curves for M dwarfs as faint as $T=15$ mag (Kunimoto et al., 2022). For TESS Sectors 61 – 65 (2023 January 18 – June 1), we obtained light curves with 200-second cadence for a total of 7,499,939 targets.
From each light curve, we removed all data points flagged as poor quality as indicated by a quality flag of 1, as well as all data within 0.2 days of the start and end of each TESS orbit, resulting in an average observational baseline per target of 25.0 days. These times correspond to high-amplitude systematics caused by scattered light from the Earth or Moon. To remove low-frequency trends caused by stellar variability, we applied the biweight detrending algorithm implemented in wotan (Hippke et al., 2019) using a detrending window of 2 days. Given that we expect most terrestrial FFP events to have typical durations of less than one day, the 2-day window was chosen to minimize the impact of event distortion caused by detrending while removing long-term astrophysical and systematic instrumental trends.
@@ 3.2 Microlensing Event Detection
Transiting exoplanets are commonly found using the BLS period search algorithm (Kovács et al., 2002), which searches for periodic dimming of a target star. The shape of this signal can be approximated as a box, i.e. the dimming initiates and terminates instantaneously with a constant flux during the dimming. BLS can also be adapted for microlensing studies by searching for non-repeating, brightening events. Though the magnification curve associated with microlensing of point-like sources is not well-approximated by a box, finite-source effects broaden this peak, resulting in light-curves that more closely resemble a box-like plateau (see §4.2 for an example displaying these effects, or Fig. 4 of Johnson et al. (2022) for further discussion). Given that the majority of TESS targets lie within $\sim 3$ kpc of Earth, terrestrial-mass FFPs would produce events with significant finite-source effects, making the use of the BLS algorithm a reasonable initial means to flag high-SNR events of interest. BLS is also computationally inexpensive, which is necessary for searching the large TESS datasets.
@@
$\sigma_{r}^{2}=\sigma_{\mathrm{bin}}^{2}-\sigma_{\mathrm{bin,exp}}^{2},$ (7)
where $\sigma_{\mathrm{bin}}$ is the weighted standard deviation of the residual light curve after binning in time with a bin-size equal to the duration of the event, and $\sigma_{\mathrm{bin,exp}}$ is the expected standard deviation of the binned light curve if the noise were uncorrelated in time. If $\sigma_{r}$ is estimated to be less than zero, we set $\sigma_{r}=0$.
In order for a signal to be considered an initial candidate to be vetted further, we required $\mathrm{SNR}>10$ and $n\geq 5$. Our requirement of $n\geq 5$ over the event duration differs slightly from other microlensing searches, which typically require a specific number of data points that are each individually above some threshold, such as a departure $3\sigma$ from the baseline (Johnson et al., 2020). Due to TESS’s rapid cadence, the number of datapoints associated with a given signal is often large, $\geq$ 20 for most potential FFP signals. By leveraging the large number of datapoints per event, signals with departures from baseline significantly below the per-datapoint uncertainty become detectable, as has been demonstrated in transit searches, which motivates our definition of detection criterion.